George Kyparisis makes bowling balls in his Miami plant. With recent increases in his​ costs, he has a newfound interest in efficiency. George is interested in determining the productivity of his organization.  He has the following data representing a month from last year and an equivalent month this​ year:  

       

Last Year

Now

Units Produced      

1,000              

1,000

Labor​ (hours)

300

250

Resin​ (pounds)

60

46

Capital Invested​ ($)

10,000

12,500

Energy​ (BTU)

2,950

2,750

a) What is the labor productivity (bowling balls per hour) last year? What is the labor productivity this year? What is the percentage change in labor productivity? (Your response as a percentage rounded to two decimal place and include a minus sign if necessary).

b) What is the resin productivity (bowling balls per pound) last year? What is the resin productivity this year? What is the percentage change in resin productivity? (Your response as a percentage rounded to two decimal places and include a minus sign if necessary).

c) What is the capital investment productivity (bowling balls per dollar) last year? What is the capital investment productivity this year? What is the percentage change in capital investment productivity? (Your response as a percentage rounded to two decimal places and include a minus sign if necessary).

d) What is the energy productivity (bowling balls per BTU) last year? What is the energy productivity this year? What is the percentage change in energy productivity? (Your response as a percentage rounded to two decimal places and include a minus sign if necessary).
Answer #1

Solution:

(a) Labor productivity:

Labor productivity = Bowling balls / Labor hours

Labor productivity (last year) = 1,000 / 300 = 3.33 bowling balls per hour

Labor productivity (this year) = 1,000 / 250 = 4 bowling balls per hour

Percentage change in labor productivity = [(labor productivity this year - labor productivity last year) / labor productivity last year] x 100

Percentage change in labor productivity = [(4 - 3.33) / 3.33] x 100

Percentage change in labor productivity = 20.12%

(b) Resin productivity:

Resin productivity = Bowling balls / Pound

Resin productivity (last year) = 1,000 / 60 = 16.67 bowling balls per pound

Resin productivity (this year) = 1,000 / 46 = 21.74 bowling balls per pound

Percentage change in resin productivity = [(resin productivity this year - resin productivity last year) / resin productivity last year] x 100

Percentage change in resin productivity = [(21.74 - 16.67) / 16.67] x 100

Percentage change in resin productivity = 30.41%

(c) Capital investment productivity:

Capital investment productivity = Bowling balls / Dollar

Capital investment productivity (last year) = 1,000 / 10,000 = 0.1 bowling balls per dollar

Capital investment productivity (this year) = 1,000 / 12,500 = 0.08 bowling balls per dollar

Percentage change in Capital investment productivity = [(Capital investment productivity this year - Capital investment productivity last year) / Capital investment productivity last year] x 100

Percentage change in Capital investment productivity = [(0.08 - 0.1) / 0.1] x 100

Percentage change in Capital investment productivity = - 20% (negative)

(d) Energy productivity:

Energy productivity = Bowling balls / BTU

Energy productivity (last year) = 1,000 / 2,950 = 0.34 bowling balls per BTU

Energy productivity (this year) = 1,000 / 2,750 = 0.36 bowling balls per BTU

Percentage change in Energy productivity = [(Energy productivity this year - Energy productivity last year) / Energy productivity last year] x 100

Percentage change in Energy productivity = [(0.36 - 0.34) / 0.34] x 100

Percentage change in Energy productivity = 5.88%